# Anonymization-Enhanced Privacy Protection for Mobile GUI Agents: Available but Invisible

### v2

@@ Abstract
Extensive experiments on the AndroidLab and PrivScreen benchmarks show that our framework substantially reduces privacy leakage across multiple models while incurring only modest utility degradation, achieving the best observed privacy–utility trade-off among existing methods. Code available at: https://github.com/one-step-beh1nd/gui_privacy_protection
@@ 3.5 Layer 3: Secure interaction proxy
Importantly, only type parameters contain textual content that may include placeholders. All spatial and gesture-based commands operate purely on numeric indices and coordinates and therefore never include anonymization tags. ÍTable 5 summarizes representative command types and their resolution logic, while Figure 4 provides a visual example of how the ’Type’ command’s argument is resolved through placeholder mapping.
@@ 4.1 Experimental Setup
For AndroidLab, we employed two cloud-based models: GPT-4o and Gemini-3-flash-preview. We compared their performance in both original (unprotected) and anonymized settings. For PrivScreen, we evaluated three multimodal models: Qwen2.5-VL-7B [2], GPT-5, and UI-TARS-7B [20]. Our method (Ours) is compared against several baselines from the DualTAP study: Original (undefended), AnyAttack [33], FOA-Attack [12], VIP [16], and DualTAP [32]. Note that the baseline results on PrivScreen are cited directly from the DualTAP paper.
@@
On AndroidLab, we employ: Success Rate (SR), Sub-Goal Success Rate (Sub-SR), Reversed Redundancy Ratio (RRR) for efficiency, and Reasonable Operation Ratio (ROR) to measure screen-change effectiveness. SR is evaluated by human experts because the original automatic evaluation metric is overly strict, often underestimating actual task completion. In contrast, Sub-SR is still automatically evaluated. On PrivScreen, utility is measured by Accuracy (Acc). Privacy leakage is assessed at the character level via Leakage Rate (LR), Match Score (MS), BLEU, and ROUGE-L. Semantic-level leakage is measured via BERTScore and Cosine Similarity (CS).
@@ 4.2 Main Results
SoM GPT-4o Original 48.55 52.23 80.71 86.69
Ours 49.28 53.26 81.26 86.80
Gemini-3-flash-preview Original 68.12 71.73 88.48 93.15
Ours 63.77 68.35 89.59 93.24
XML GPT-4o Original 45.65 48.78 102.16 90.72
Ours 42.34 46.12 103.35 91.54
Gemini-3-flash-preview Original 48.91 53.45 107.25 98.55
Ours 52.17 57.17 98.91 95.68
@@
The main results on the AndroidLab benchmark are reported in Table 7. Overall, introducing our privacy protection layer leads to a small and controlled reduction in task success, while preserving the rationality and stability of agent behaviors. Under the SoM agent mode, applying our method slightly decreases the SR and Sub-SR for both backbone models. For GPT-4o, SR (human-evaluated) drops from 36.23% to 32.61%, while Sub-SR (automatically evaluated) shows a similar mild decrease. However, the ROR remains stable (e.g., 86.69% to 86.80% for GPT-4o), indicating that our protection does not introduce irrational interactions. Under XML mode, the impact is even milder; for Gemini-3-flash-preview, our method even improves SR from 34.78% to 39.86%, suggesting that anonymization may reduce distractions and help the agent focus on task-relevant structures.
@@
Table 8 presents the results on the PrivScreen benchmark. Our method consistently achieves the strongest privacy protection across all models. For Qwen2.5-VL-7B, our method maintains the same task accuracy (89.00%) while dramatically reducing the LR from 97.14% to 19.52% and MS from 96.99% to 30.69%. For GPT-5, we preserve high utility (90.00% vs. 93.00%) while reducing BertScore from 0.9260 to 0.1495. Compared to existing defenses such as AnyAttack or DualTAP, our method consistently achieves the lowest leakage-related metrics while maintaining competitive accuracy.
@@ 5 Conclusion
Extensive experiments on the AndroidLab and PrivScreen benchmarks demonstrate that our approach consistently achieves the lowest observed privacy leakage across multiple models, while incurring only modest utility degradation. These results indicate a strong and favorable privacy–utility trade-off compared to existing methods, validating the practicality of our framework for real-world mobile GUI agent scenarios.

### v3

@@ Abstract
Extensive experiments on the AndroidLab and PrivScreen benchmarks show that our framework substantially reduces privacy leakage across multiple models while incurring only modest utility degradation, yielding a favorable privacy–utility trade-off compared with existing methods. Code available at: https://github.com/one-step-beh1nd/gui_privacy_protection
@@ 3.5 Layer 3: Secure interaction proxy
Importantly, only type parameters contain textual content that may include placeholders. All spatial and gesture-based commands operate purely on numeric indices and coordinates and therefore never include anonymization tags. Table 5 summarizes representative command types and their resolution logic, while Figure 4 provides a visual example of how the ’Type’ command’s argument is resolved through placeholder mapping.
@@ 4.1 Experimental Setup
For AndroidLab, we employed two cloud-based models: GPT-4o and Gemini-3-flash-preview. Besides the unprotected input (Original) and our anonymization pipeline (Ours), we include a strong privacy baseline, Full cover, which replaces all detected sensitive spans with a fixed token ([privacy information]) instead of category-preserving placeholders. In SoM mode, we further include the screenshot perturbation method from DualTAP [32] as an additional baseline (DualTAP); this baseline is not applicable to XML mode because XML mode does not consume screenshot inputs. For PrivScreen, we evaluated three multimodal models: Qwen2.5-VL-7B [2], GPT-5, and UI-TARS-7B [20]. Our method (Ours) is compared against several baselines from the DualTAP study: Original (undefended), AnyAttack [33], FOA-Attack [12], VIP [16], and DualTAP [32]. Note that the baseline results on PrivScreen are cited directly from the DualTAP paper.
@@
On AndroidLab, we employ: Success Rate (SR), Sub-Goal Success Rate (Sub-SR), Reversed Redundancy Ratio (RRR) for efficiency, and Reasonable Operation Ratio (ROR) to measure screen-change effectiveness. The evaluation protocol for SR depends on task type. For operation tasks, we arrange all execution screenshots in chronological order and provide them with the task description to a vision model, which judges whether the task is successfully completed; the vision evaluator is Qwen3-VL-235B-A22B-Thinking. For query_detect tasks, we provide the task description, reference answer, and the GUI agent’s answer to a text model, which judges whether the response is correct; the text evaluator is DeepSeek-V3.2-Instruct. Sub-SR, RRR, and ROR are still computed with AndroidLab’s original automatic evaluation pipeline. On PrivScreen, utility is measured by Accuracy (Acc). Privacy leakage is assessed at the character level via Leakage Rate (LR), Match Score (MS), BLEU, and ROUGE-L. Semantic-level leakage is measured via BERTScore and Cosine Similarity (CS).
@@ 4.2 Main Results
SoM GPT-4o Original 44.20 81.08 90.00 89.34
Full cover 34.78 76.18 86.16 87.48
DualTAP 33.33 71.65 72.16 85.77
Ours 40.58 79.43 80.47 86.68
Gemini-3-flash-preview Original 70.29 88.45 87.50 94.43
Full cover 55.80 82.84 85.75 92.32
DualTAP 63.77 85.59 82.62 92.72
Ours 63.77 85.43 87.04 93.73
XML GPT-4o Original 35.51 74.30 92.40 90.12
Full cover 39.86 76.07 93.62 90.84
Ours 41.30 77.36 92.15 89.55
Gemini-3-flash-preview Original 57.97 85.28 91.86 93.73
Full cover 50.00 83.49 94.07 92.85
Ours 54.34 84.32 90.91 92.67
@@
The main results on the AndroidLab benchmark are reported in Table 7. Overall, our method introduces only limited utility degradation after privacy protection, and in most settings clearly outperforms Full cover and DualTAP, showing that category-preserving anonymization provides a favorable privacy–utility trade-off.
Under SoM mode, the unprotected setting achieves the highest SR and Sub-SR for both backbones, but our method remains notably closer to Original than the two baselines. For GPT-4o, Original reaches 44.20% SR, while Full cover and DualTAP drop to 34.78% and 33.33%; our method still achieves 40.58%, with Sub-SR also higher than both baselines. For Gemini-3-flash-preview, our SR is 63.77%, matching DualTAP and clearly above Full cover (55.80%), while maintaining a high ROR of 93.73%. This suggests that our method preserves more stable interaction behavior under multimodal inputs.
Under XML mode, the methods exhibit a finer-grained trade-off. For GPT-4o, our method achieves the best SR and Sub-SR (41.30%/77.36%), outperforming both Original and Full cover, while Full cover remains best on RRR and ROR. For Gemini-3-flash-preview, Original is still strongest on SR, Sub-SR, and ROR, and Full cover is best on RRR; our method lies between them overall, retaining strong task completion while avoiding the information loss caused by blanket replacement.
@@
Table 8 presents the results on the PrivScreen benchmark. Overall, our method provides strong privacy protection across all tested models while keeping utility competitive. For Qwen2.5-VL-7B, our method maintains the same task accuracy (89.00%) while dramatically reducing the LR from 97.14% to 19.52% and MS from 96.99% to 30.69%. For GPT-5, we preserve high utility (90.00% vs. 93.00%) while reducing BertScore from 0.9260 to 0.1495. Compared to existing defenses such as AnyAttack or DualTAP, our method consistently improves key leakage-related metrics while maintaining competitive accuracy.
@@ 5 Conclusion
Extensive experiments on the AndroidLab and PrivScreen benchmarks demonstrate that our approach consistently reduces privacy leakage across multiple models, while incurring only modest utility degradation. These results indicate a strong and favorable privacy–utility trade-off compared to existing methods, validating the practicality of our framework for real-world mobile GUI agent scenarios.